40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Apr 6 02:07:40 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
    Reply  Wed Apr 10 00:33:09 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source Modulation_frequency_100Hz.jpgModulation_frequency_30Hz.jpgModified_setup.JPG
       Reply  Wed Apr 10 16:58:54 2019, rana, Update, IOO, fiber MZ for NPRO freq noise measurements 
          Reply  Wed Apr 10 22:59:22 2019, gautam, Update, IOO, Spooled fiber 
             Reply  Thu Apr 11 09:05:06 2019, Anjali, Update, IOO, Spooled fiber Frequecy_modulation_32_Hz.pdfFrequecy_modulation_140_Hz.pdfFrequecy_modulation_300_Hz.pdfWithout_modulation.pdfNew_fiber_spool.JPG
       Reply  Fri Apr 12 01:22:27 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
          Reply  Mon Apr 15 22:39:10 2019, gautam, Update, Frequency noise measurement, Alternate setup with PSL pickoff 
             Reply  Wed Apr 17 00:43:38 2019, gautam, Update, Frequency noise measurement, MZ interferometer ---> DAQ 
                Reply  Thu Apr 25 03:32:25 2019, Anjali, Update, Frequency noise measurement, MZ interferometer ---> DAQ Time_domain_output.pdfFrequency_noise.pdfschematic_heterodyne_setup.pngNoise_budget_1_micron_in_Hz_per_rtHz.pdf
                   Reply  Thu Apr 25 10:25:19 2019, gautam, Update, Frequency noise measurement, Homodyne v Heterodyne 
                      Reply  Thu Apr 25 15:47:54 2019, Anjali, Update, Frequency noise measurement, Homodyne v Heterodyne Experimental_setup.JPG
Message ID: 14571     Entry time: Thu Apr 25 03:32:25 2019     In reply to: 14547     Reply to this: 14573
Author: Anjali 
Type: Update 
Category: Frequency noise measurement 
Subject: MZ interferometer ---> DAQ 
  • Attachment #1 shows the time domain output from this measurement. The contrast between the maximum and minimum is better in this case compared to the previous trials.
  • We also tried to extract the frequency noise of the laser from this measurement. Attachment #2 shows the frequency noise spectrum. The experimental result is compared with the theoretical value of frequency noise. Above 10 Hz, the trend is comparable to the expected 1/f characteristics, but there are other peak also appearing. Similarly, below 10 Hz, the experimentally observed value is higher compared to the theory.
  • One of the uncertainties in this result is because of the length fluctuation of the fiber. The phase fluctuation in the system could be either because of the frequency noise of the laser or because of the length fluctuation of the fiber.  So,one of the reasons for the discrepancy between the experimental result and theory could be because of  fiber length fluctuation. Also, there were no locking method been applied to operate the MZI in the linear range.
  • The next step would be to do a heterodyne measurement. Attachment #3 shows the schematic for the heterodyne measurement. A free space AOM can be inserted in one of the arms to do the frequency shift. At the output of photodiode, a RF heterodyne method as shown in attachment #3 can be applied to separate the inphase and quadrature component. These components need to be saved with a deep memory system. Then the phase and thus the frequency noise can be extracted.
  • Attachment #4 shows the noise budget prepared for the heterodyne setup. The length of the fiber considered is 60 m and the photodiode is PDA255. I also have to add the frequency noise of the RF driver and the intensity noise of the laser in the noise budget.
Quote:
  1. Delay fiber was replaced with 5m (~30 nsec delay)
    • The fringing of the MZ was way too large even with the free running NPRO (~3 fringes / sec)
    • Since the V/Hz is proportional to the delay, I borrowed a 5m patch cable from Andrew/ATF lab, wrapped it around a spool, and hooked it up to the setup
    • Much more satisfactory fringing rate (~1 wrap every 20 sec) was observed with no control to the NPRO
  2. MZ readout PDs hooked up to ALS channels
    • To facilitate further quantitative study, I hooked up the two PDs monitoring the two ports of the MZ to the channels normally used for ALS X.
    • ZHL3-A amps inputs were disconnected and were turned off. Then cables to their outputs were highjacked to pipe the DC PD signals to the 1Y3 rack
    • Unfortunately there isn't a DQ-ed fast version of this data (would require a model restart of c1lsc which can be tricky), but we can already infer the low freq fringing rate from overnight EPICS data and also use short segments of 16k data downloaded "live" for the frequency noise measurement.
    • Channels are C1:ALS-BEATX_FINE_I_IN1 and C1:ALS-BEATX_FINE_Q_IN1 for 16k data, and C1:ALS-BEATX_FINE_I_INMON and C1:ALS-BEATX_FINE_I_INMON for 16 Hz.

At some point I'd like to reclaim this setup for ALS, but meantime, Anjali can work on characterization/noise budgeting. Since we have some CDS signals, we can even think of temperature control of the NPRO using pythonPID to keep the fringe in the linear regime for an extended period of time.

Attachment 1: Time_domain_output.pdf  31 kB  Uploaded Thu Apr 25 04:33:01 2019  | Hide | Hide all
Time_domain_output.pdf
Attachment 2: Frequency_noise.pdf  42 kB  Uploaded Thu Apr 25 04:33:43 2019  | Hide | Hide all
Frequency_noise.pdf
Attachment 3: schematic_heterodyne_setup.png  96 kB  Uploaded Thu Apr 25 04:33:55 2019  | Hide | Hide all
schematic_heterodyne_setup.png
Attachment 4: Noise_budget_1_micron_in_Hz_per_rtHz.pdf  45 kB  Uploaded Thu Apr 25 04:34:19 2019  | Hide | Hide all
Noise_budget_1_micron_in_Hz_per_rtHz.pdf
ELOG V3.1.3-